# CroSTAta: Cross-State Transition Attention Transformer for Robotic Manipulation

### v1

@@ IV-C Baseline Methods
We compare our STA Transformer against four baseline approaches: a Transformer that processes temporal sequences with standard cross-attention, one that does not take past historical context into account, and established temporal modeling approaches including TCN [27] and LSTM [28] networks. All networks present relatively small architectures with 4 processing layers and a hidden size of 512. We equipped TCN and LSTM networks with additional final feed-forward networks to predilige a fair performance comparison given the parameter count differences with transformer-based methods. The encoder structures remain comparable across all methods, with a hidden representation of 512. Additional implementation details and training hyperparameters are provided in Table I.
@@
IV-D Performance Evaluation
Fig. 5 presents the performance results in terms of success rate, demonstrating the relative effectiveness of different temporal modeling approaches when learning from recovery-rich demonstrations. Our STA Transformer consistently outperforms all baseline approaches across all four tasks, with particularly notable improvements in precision-requiring tasks such as PegInsertionSide, where STA achieves 18.3% success compared to 7.7% for standard Transformer – representing more than a 2x improvement. Interestingly, the standard Transformer approach with full history attention show occasional underperformance compared to no-history baselines in some tasks, suggesting that naive historical information integration does not always benefit policy performance. Additionally, traditional sequence modeling approaches demonstrate notable limitations in this domain, with TCN and LSTM networks consistently showing lower performance. This is particularly true for LSTM, which performs poorly across precision-critical tasks. These results support our hypothesis that superior temporal reasoning capabilities, achieved through our structured attention mechanisms designed for state transition modeling, provide performance advantages compared to both history-agnostic approaches and standard sequence modeling techniques.
Figure 6: Attention pattern analysis during a PegInsertionSide trajectory showing four critical execution phases. Top heatmaps display attention weights averaged across heads for state tokens over a 15-step history window (darker = higher attention). Bottom graphs show per-head attention values summed across tokens. In both visualizations, timestep 0 represents the oldest historical information while timestep 15 represents the present. Vertical dotted lines indicate the corresponding timestep in the temporal sequence for each colored execution phase. The visualization demonstrates how STA learns to retrieve relevant historical context during recovery phases (t=29) compared to initial execution patterns (t=4, t=10).
@@ IV-E Attention Pattern Analysis
Beyond overall performance metrics, we analyze the learned attention patterns to understand how our proposed mechanism processes temporal information. Specifically, we examine how STA weights historical attention scores, e.g., $Q_{t-k}K^{T}_{t-k}$ relative to $k$ timestep in the past, through the state transition scores from $S_{t-k}S^{T}_{t}$ when situations requiring corrective actions occur. Fig. 6 shows a trajectory executed by our transformer-based policy in the PegInsertionSide task, which involves reaching the peg, lifting it from the white end, and inserting it into the box hole. The analyzed scenario demonstrates four key phases: starting phase (t=4), failed grasping attempt (t=10), successful recovery grasp (t=29), and final insertion approach (t=39). We examine the patterns of attention modules at these critical moments by observing how the value of state transition scores evolves over time. The visualization shows the values extracted from the last Transformer layer, which we observed to provide the most interpretable results. The top heatmaps display state transition weights averaged across all heads for state tokens, showing their individual relevance across the 15-step history window (darker colors indicate stronger attention). The bottom graphs show state transition values summed across tokens per computation head, revealing which heads contribute the most to historical context retrieval at each timestep. Starting from a pre-initialized history of identical starting states – strategy adopted strictly for analysis purposes – we observe initial patterns showing lower state transition scores for past states and higher scores for current state information (t=4). This pattern persists through the first grasping attempt (t=10) that ultimately fails. The comparison between the first and second grasping attempts reveals a striking change: during the successful recovery attempt (t=29), the patterns show higher state transition scores extending more into the past, with particular activation of heads 2 and 4 that appear to function as retrieval pathways for relevant historical state relationships. The final trajectory phase (t=39) shows restored focus on recent timesteps with higher state transition scores on specific tokens, likely facilitating the precise movements necessary for successful peg insertion.
This analysis provides insights into how the model attends to relevant past events in its execution history during corrective reattempts, while downweighting attention scores in other scenarios, supporting our hypothesis that structured attention mechanisms can better exploit the temporal dependencies present in recovery-rich demonstrations.
@@ IV-F Impact of Temporal Masking on Training and Inference Robustness
Figure 7: Ablation study comparing temporal masking effects during training and robustness at inference time. Results show the StackCube task performance under different training (standard vs. masked) and inference (complete vs. partially masked observations) conditions. Our proposed Transformer with STA benefits from temporal masking training while standard Transformer shows minimal improvement, demonstrating architecture-specific advantages.
To understand the specific contributions of our temporal masking training strategy, we evaluate both the direct effects of masked training and the robustness it provides at inference time. We compare standard training with masked training under inference conditions with both complete and partially masked observations. Specifically, for ’masked’ inference evaluation, we remove visual observations for $n$ consecutive steps where $n$ ranges from 0 to LL-1, with $L$ being the total sequence length. ensuring at least one timestep retains the full observation. The masking duration $n$ is resampled with 0.1 probability when it reaches zero otherwise standard conditions of inference apply. Fig. 7 presents results for the StackCube task, revealing several important findings. Our STA Transformer trained with masking achieves superior performance under standard inference conditions compared to the same model trained without masking (71.3% vs 64.7%), demonstrating that the temporal masking strategy enhances learning even when full observations are available during deployment. This suggests that forcing the model to rely on historical context during training develops more robust temporal reasoning capabilities. In contrast, the standard Transformer baseline does not benefit from masked training under normal inference conditions, with performance remaining essentially unchanged compared to standard training (65.0% vs 62.0%). This indicates that the effectiveness of temporal masking is specifically tied to our state transition attention mechanism rather than being a general improvement applicable to any architecture. Notably, when evaluating robustness under masked inference conditions, our STA Transformer trained with masking strategy maintains a significant advantage over the baseline (52.3% vs 42.3% under masked inference).
@@ IV-G Historical Context Dependency
We provide additional analysis of how history length affects the performance of our STA Transformer. We focus on StackCube and PegInsertionSide environments, training our method with 15 historical steps (16-step sequences) and then evaluating performance with variable history lengths. Additionally, we train reference policies on sequences with history lengths of 7, 3, 1, and 0 steps to provide baseline performance where training and inference settings match. Results in Fig. 8 demonstrate the robustness of our approach when evaluated with limited historical information available. For PegInsertionSide, we observe only modest performance degradation as inference history decreases, while in StackCube, our method occasionally even exceeds the initial performance. This robustness suggests that our STA mechanism enables learning effective decision-making from rich historical information during training, which transfers well to inference scenarios with truncated historical context. Relative to reference policies trained with shorter histories, we observe substantial performance drops, particularly for those trained with 1 and 3 historical steps. We attribute this degradation to our temporal masking procedure during training, which may excessively dampen the training signal when applied to sequences with already limited temporal information. Notably, the reference policy with 0 historical steps at training (where no temporal masking is applied), presents higher performance than those with 1 and 3 historical steps. This interaction between masking strategy and history length highlights the importance of adequate temporal context for our training approach to be effective.
Figure 8: Historical context dependency analysis showing performance vs. inference history length for Transformers with STA trained with 15-step history. Star markers indicate reference performance from models trained with matching history lengths. Results demonstrates robust performance across different inference history lengths, with StackCube showing remarkable stability and PegInsertionSide exhibiting graceful degradation.
@@ V Discussion
Our work demonstrates a key insight on how our proposed structured attention mechanisms can effectively learn to exploit temporal dependencies present in recovery-rich demonstrations. This is evidenced by attention pattern analysis showing adaptive retrieval of historical information during recovery phases. However, we acknowledge limitations constraining the generalizability of our findings. The manipulation tasks evaluated are relatively short-horizon and do not strictly necessitate historical reasoning to succeed, as evidenced by the non-trivial performance of baselines omitting historical context. While our improvements target challenging cases where imprecise movements and unrecoverable situations occur, more complex temporally-extended tasks would provide stronger evaluation grounds for historical reasoning capabilities. Additionally, our data collection methodology remains constrained by the capabilities of privileged policies used to generate failure and recovery patterns, limiting the diversity and complexity of temporal dependencies that can be learned.
@@ VI Conclusions
We have presented State Transition Attention (STA), a novel attention mechanism that modulates attention weights based on learned state evolution patterns to improve temporal reasoning in robotic manipulation policies. Our experimental evaluation demonstrates that STA consistently outperforms standard temporal modeling approaches across four ManiSkill manipulation tasks, with particularly notable improvements in precision-critical scenarios. The ablation studies reveal additional findings regarding temporal masking and demonstrate robustness to reduced historical context at inference time. These results establish our approach as a promising direction for developing more capable manipulation policies that can effectively learn from and reason about their execution history. Future work should explore more sophisticated data collection strategies and evaluate performance on complex temporally-extended tasks that more strongly necessitate historical reasoning capabilities.

### v2

@@ IV-C Baseline Methods
We compare our STA Transformer against five baselines: a Transformer with standard cross-attention over state sequences, a Transformer that does not take past historical context into account, a self-attention-only Transformer processing state and input tokens together, and established temporal modeling approaches including TCN [17] and LSTM [11] networks. All networks use 4 layers and hidden size 512. We equipped TCN and LSTM networks with additional final feed-forward networks to predilige a fair performance comparison given the parameter count differences with transformer-based methods. The encoder structures remain comparable across all methods, with a hidden representation of 512. Additional implementation details and training hyperparameters are provided in Table I.
IV-D Performance Evaluation
Fig. 5 presents the performance results in terms of success rate, demonstrating the relative effectiveness of different temporal modeling approaches when learning from recovery-rich demonstrations. Our STA Transformer outperforms all baselines across the three precision-requiring and coordination-demanding tasks, with particularly notable improvements on PegInsertionSide, reporting more than a 2× improvement over standard Transformer (18.3% vs 7.7%). Notably, the Transformer with self-attention only performs similarly to or below its cross-attention variant, while both acting with full history occasionally underperform the no-history baseline, together suggesting that naive integration of historical information does not always benefit policy performance. Traditional sequence modeling approaches show consistent limitations in this domain, with LSTM performing particularly poorly on precision-critical tasks. On UnitreeG1TransportBox, performance is comparably high across all methods. We attribute this to the task’s inherent robustness to the noise introduced during data collection: the privileged policy used for demonstration generation recovers reliably without exhibiting meaningful corrective behaviors, limiting the temporal structure present in the training data and consequently reducing the signal available for STA to exploit. These results support our hypothesis that structured attention mechanisms designed for state transition modeling can provide meaningful advantages over history-agnostic and standard sequence modeling approaches, with gains amplified where demonstrations contain rich temporal structure and tasks present challenging execution conditions.
Figure 6: Attention pattern analysis during a PegInsertionSide trajectory showing four critical execution phases. Top heatmaps display state transition weights averaged across heads for individual state tokens over a 15-step history window (darker = higher attention). Bottom graphs show the complementary view: per-head attention values summed across all tokens, revealing which heads contribute most to historical retrieval. Both visualizations share the same temporal axis where timestep 0 represents the oldest historical information and timestep 15 the present. Vertical dotted lines indicate the corresponding timestep in the temporal sequence for each colored execution phase. The visualization demonstrates how STA learns to selectively retrieve relevant historical context during recovery phases (t=29) compared to initial execution patterns (t=4, t=10).
@@ IV-D Performance Evaluation
Parameters 9.7M 10.4M 21.2M (Transformer)
17M (Transformer SA only)
23.3M (Transformer w/ STA)
@@ IV-E Attention Pattern Analysis
Beyond overall performance metrics, we analyze the learned attention patterns to understand how STA processes temporal information. We examine state transition scores $S_{t-k:t}S^{T}_{t}$ that modulate historical attention scores $Q_{t-k:t}K^{T}_{t-k:t}$ at timestep $t$. Fig. 6 shows a trajectory executed by our transformer-based policy in the PegInsertionSide task, which involves reaching the peg, lifting it from the white end, and inserting it into the box hole. We analyze how state transition values evolve across four key execution phases: starting phase (t=4), failed grasping attempt (t=10), successful recovery grasp (t=29), and final insertion approach (t=39). The visualization shows values extracted from the last Transformer layer, which provides the most interpretable results. The top heatmaps display state transition weights averaged across all heads for state tokens, showing their individual relevance across the 15-step history window (darker colors indicate stronger attention). The bottom graphs show state transition values summed across tokens per computation head, revealing which heads contribute the most to historical context retrieval at each timestep. Starting from a pre-initialized history of identical starting states – strategy adopted strictly for analysis purposes – we observe initial patterns showing lower state transition scores for past states and higher scores for current state information (t=4). This pattern persists through the first grasping attempt (t=10) that ultimately fails. The comparison between the first and second grasping attempts reveals a striking change: during the successful recovery attempt (t=29), the patterns show higher state transition scores extending more into the past, with particular activation of heads 2 and 4 that appear to function as retrieval pathways for relevant historical state relationships. The final trajectory phase (t=39) shows restored focus on recent timesteps with higher state transition scores on specific tokens, likely facilitating the precise movements necessary for successful peg insertion.
This analysis provides insights into how STA learns to selectively attend to relevant past events during challenging execution phases, while downweighting irrelevant historical information in others, supporting our hypothesis that structured attention mechanisms can better exploit the temporal dependencies present in recovery-rich demonstrations.
@@ IV-F Impact of Temporal Masking on Training and Inference Robustness
Figure 7: Ablation study comparing temporal masking effects during training and robustness at inference time. Results show the StackCube task performance under different training (standard vs. masked) and inference (complete vs. partially masked observations) conditions. Differently from baseline approaches our proposed STA Transformer benefits from temporal masking training, demonstrating architecture-specific advantages.
To understand the specific contributions of our temporal masking training strategy, we evaluate both its direct effects on performance and the robustness it provides at inference time. We compare policies from standard and masked training at inference with both complete and partially masked observations. For ’masked’ inference evaluation, we remove visual observations for $n$ consecutive steps where $n$ ranges from 0 to LL-1, with $L$ being the total sequence length, ensuring at least one timestep retains the full observation, with n resampled with 0.1 probability when it reaches zero, otherwise standard inference conditions apply. Fig. 7 presents results for the StackCube task, revealing several important findings. Our STA Transformer trained with masking achieves superior performance under standard inference conditions compared to the same model trained without masking (71.3% vs 64.7%), demonstrating that temporal masking enhances learning even when full observations are available during deployment. This suggests that forcing the model to rely on historical context during training develops more robust temporal reasoning capabilities. In contrast, neither of the other Transformer baselines benefits from masked training – the standard Transformer loses 3.0% and the SA only variant 4.4% – confirming that the effectiveness of temporal masking is specifically tied to our state transition attention mechanism rather than being a general improvement applicable to any architecture. Notably, under masked inference conditions, our STA Transformer maintains a significant advantage over baselines (52.3% vs 42.3% and 37.7%).
@@ IV-G Historical Context Dependency
We analyze how history length at inference affects the performance of our STA Transformer, focusing on StackCube and PegInsertionSide. We train our method with 15 historical steps (16-step sequences) and evaluate with variable history lengths. Additionally, we train reference policies – STA Transformers trained with history lengths of 7, 3, 1, and 0 steps – to provide baseline performance where training and inference settings match. Results in Fig. 8 demonstrate the robustness of our approach when evaluated with limited historical information available. For PegInsertionSide, we observe only modest performance degradation as inference history decreases, while in StackCube, our method occasionally even exceeds the initial performance. This robustness suggests that our STA mechanism enables learning effective decision-making from rich historical information during training, which transfers well to inference scenarios with truncated historical context. Relative to reference policies trained with shorter histories, we observe substantial performance drops, particularly for those trained with 1 and 3 historical steps. We attribute this degradation to our temporal masking procedure during training, which may excessively dampen the training signal when applied to sequences with already limited temporal information. Notably, the trained reference policy with 0 historical steps at training (where no temporal masking is applied), presents higher performance than those with 1 and 3 historical steps. This interaction between masking strategy and history length highlights the importance of adequate temporal context for our training approach to be effective.
Figure 8: Historical context dependency analysis showing performance vs. inference history length for a STA Transformer trained with 15-step history. Star markers indicate reference performance from models trained with inference-matching history lengths. Results demonstrates robust performance across different inference history lengths, with StackCube showing remarkable stability and PegInsertionSide exhibiting graceful degradation.
@@ V Discussion
Our work demonstrates how structured attention mechanisms can effectively exploit temporal dependencies present in recovery-rich demonstrations, as evidenced by attention pattern analysis showing adaptive retrieval of historical information during recovery phases and by consistent performance improvements across precision-requiring and coordination-demanding tasks. STA’s advantage emerges specifically in scenarios combining rich temporal structure in training data with challenging execution conditions, while benefits diminish when demonstrations present limited diversity in failure and recovery patterns, as observed in UnitreeG1TransportBox. We acknowledge limitations constraining the generalizability of our findings. The evaluated tasks are relatively short-horizon and, as evidenced by non-trivial baseline performance without historical context, do not fundamentally require temporal reasoning for basic task completion. Our improvements specifically target the challenging cases where imprecise movements or unrecoverable situations benefit from historical awareness. More complex temporally-extended tasks with stronger partial observability – where current observations alone would be insufficient for informed decision-making – would provide additional validation of historical reasoning capabilities, though extending to such scenarios is primarily limited by hardware requirements for training with extended sequences and storing larger histories at inference. Future extensions should address scalability concerns through memory-efficient techniques for both training and inference phases. Our evaluation is conducted entirely in simulation, though no inherent architectural barriers prevent real-world deployment beyond standard sim-to-real transfer challenges common to vision-based policies. Additionally, our data collection methodology remains constrained by the capabilities of privileged policies used to generate failure and recovery patterns, limiting the diversity and complexity of temporal dependencies that can be learned. Alternative approaches including human demonstrations of natural recovery behaviors or online learning could provide richer temporal structure for future investigation.
@@ VI Conclusions
We have presented CroSTAta, employing a State Transition Attention mechanism that modulates attention weights based on learned state evolution patterns to improve temporal reasoning in robotic manipulation policies. Our experimental evaluation demonstrates that STA outperforms standard temporal modeling approaches across manipulation tasks, with particularly notable improvements in precision-critical scenarios. The ablation studies reveal additional findings regarding temporal masking and demonstrate robustness to reduced historical context at inference time. These results establish our approach as a promising direction for developing more capable manipulation policies that can effectively learn from and reason about information history.